OpenStudy (anonymous):

what the solution of the equation? x x-4 - = - 6 2

OpenStudy (zepp):

Cross something theorem 2 * x = (x-4) * 6

OpenStudy (zepp):

2x = 6 ( x-4) 2x = 6x -24 0 = 6x -24- 2x 0 = 4x - 24 24 = 4x x = 6